OpenStudy (clara1223):

how to find the limit as x approaches 4 of (4-x)/(2-sqrt(x))

OpenStudy (freckles):

hint: \[4-x=(2-\sqrt{x})(2+\sqrt{x})\]

OpenStudy (clara1223):

\[\lim_{x \rightarrow 4}\frac{ 4-x }{ 2-\sqrt{x} }\]

OpenStudy (freckles):

are you cool?

OpenStudy (freckles):

like with the hint

OpenStudy (clara1223):

Not really haha I remember learning that in class but I don't remember how to simplify from there.

OpenStudy (freckles):

so you never canceled common factors in division ?

OpenStudy (freckles):

for example a/a=1 do you know this?

OpenStudy (freckles):

\[\lim_{x \rightarrow 4} \frac{(2-\sqrt{x})(2+\sqrt{x})}{2-\sqrt{x}}\] you should see something to cancel here

OpenStudy (dinamix):

4 this lim x--->4

OpenStudy (clara1223):

@freckles yes i do. do i plug 4 into the remaining equation?

OpenStudy (freckles):

yeah